Are the two line parallel, perpendicular, or neither?
16x+2y=18; 3y=-24x+45
Show work.

Answer by mananth(16949)

16x+2y=18;
2y= 18-16x
/2
y = 9-8x
y=-8x+9.................(1)

3y=-24x+45
/3
y= -8x+15...............(2)
compare (1) & (2)
slope m= -8 in both so the lines are parallel